OpenStudy (anonymous):

Find the first four terms of the given sequence tn = n/(n+1). A. ½, 2/3, ¾, 4/5 B. 1, ½, 1/3, ¼ C. 1, ½, 2/3, ¾ D. 1, 2, 3, 4

jimthompson5910 (jim_thompson5910):

when n = 1, what is tn?

jimthompson5910 (jim_thompson5910):

tn = n/(n+1) t1 = 1/(1+1) ... plug in n = 1 t1 = ???